OpenStudy (anonymous):

compute difference quotient of -1/(x-5)

OpenStudy (anonymous):

im supposed to use f(a+h)-f(a)/h

OpenStudy (anonymous):

any ideas? you have to compute \[\frac{\frac{-1}{a+h-5}+\frac{1}{a-5}}{h}\]

OpenStudy (anonymous):

lets forget the h for a moment and concentrate on the numerator, because that is where all the algebra is

OpenStudy (anonymous):

\[\frac{1}{a-5}-\frac{1}{a+h-5}\] \[\frac{a+h-5-(a-5)}{(a-5)(a+h-5)}\] \[\frac{h}{(a-5)(a+h-5)}\]

OpenStudy (anonymous):

now divide by h, and of course the h cancels, leaving \[\frac{1}{(a-5)(a+h-5)}\]